Dearborn Supplies has total sales of

$ 207

​million, assets of

$ 108

​million, a return on equity of

35

​percent, and a net profit margin of

7.9

percent. What is the​ firm's debt​ ratio?

The​ company's debt ratio is

nothing​%.

​(Round to one decimal​ place.)
Answer #1

Profit margin=Net income/Sales

Net income=(207*7.9%)=$16.353 million

ROE=Net income/equity

equity=16.353/0.35=$46.7228571 million

Total assets=debt+equity

debt=(108-46.7228571)=$61.2771429million

Hence debt ratio=debt/Total assets

=61.2771429/108

=56.7%(Approx).